You can calculate how far the moon is away ( assuming yu set your laser to stun , so yu dont blow up the moon ) the speed of light is 3 x 10[sup]8[/sup] m/s . When the moon is at apogee , the laser would take 2.7 seconds for it to leave earth , reflect off the moon , and return back to earth. knowing this , calculate the distance to the moon. ( hint* - your answer should be in meters )
1
Expert's answer
2011-09-18T16:13:58-0400
We are given:
Speed of light
c=3*108ms
Time for light to leave earth, reflect off the moon,and return back to earth:
t=2.7 s
So light traveleddistance from Erath to Moon twice ( to Moon and back to Earth)<br/>Distance isvelocity multiplied by time so:
2*L=c*t

Answer: 4.05*108 m